OpenStudy (anonymous):

log3 x + log3(x-24)=4

OpenStudy (anonymous):

ok i think i can answer this one... but i dont know if its the equation that you need, but let me try.

OpenStudy (anonymous):

3x+3(x-24)=4 3x+3(x-24+24)=4+24 3x+3x=28 6x=28 6x/6x=28/6 x=4.6 and over the six put a repeating bar. :)

OpenStudy (anonymous):

is that how you needed it? i didnt understand the log part.... im only in sixth grade taking seventh grade math, but we are learning two step equations, three step equations, and equations with variables on both sides.

OpenStudy (dumbcow):

hey that was pretty good for sixth grade ;) yes the log part changes the answer a bit and you have to multiply the 24 by 3 before you add it to the other side. good job though keep it up

OpenStudy (dumbcow):

ok for this problem you need to combine the logs into one log using the rule logx + logy = log(xy) => log(x^2-24x) = 4 Since the base is 3 i can say 3 raised to both sides canceling out the log => x^2 - 24x = 3^4 = 81 set equal to 0 =>x^2 - 24x - 81 = 0 Factor and solve for x. **Note x cannot be negative because you cant take log of neg number Hope this help